OpenStudy (anonymous):

Identify the solution set of the inequality using the given replacement set. x < –7; {–15, –10.5, –7.2, –7, 0, 6} A. {–15, –10.5, –7.2, –7} B. {–15, –10.5, –7.2} C. {–15, –10.5, 0, 6} D. {0, 6}

OpenStudy (anonymous):

a?

OpenStudy (johnp12):

I think it maybe (a) but not to sure hold on a sec

OpenStudy (freckles):

is -7 less than -7?

OpenStudy (freckles):

-7=-7 so -7 is neither less than -7 nor greater than do you happen to have an equal sign also since you are including -7 in your solution?

OpenStudy (freckles):

\[x \le -7 \text{ or } x<-7?\] \[x \le -7 \text{ says all x less than or equal to -7 }\] \[x < -7 \text{ says all x less than -7 }\]

OpenStudy (anonymous):

im confused

OpenStudy (freckles):

your inequality says x<-7 is -7<-7?

OpenStudy (freckles):

I'm asking since you including -7 in your solution?

OpenStudy (anonymous):

no its equal

OpenStudy (freckles):

so -7 would definitely not be a solution of x<-7 since it does not satisfy the inequality \[\text{ unless you meant } x \le -7\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

@chosenmatt

OpenStudy (freckles):

You are still not convinced -7 should not be in the solution set?

OpenStudy (freckles):

-7 equals -7 it is definitely not less than -7... so all the other numbers in the set were less than -7

OpenStudy (freckles):

which set includes all the numbers you said before without the -7

OpenStudy (anonymous):

a

OpenStudy (freckles):

no set a includes -7

OpenStudy (freckles):

we already said -7 is not a solution of the inequality x<-7

OpenStudy (anonymous):

b

OpenStudy (anonymous):

no c

OpenStudy (freckles):

yes all the numbers in b are in a except the -7

OpenStudy (freckles):

and no not c

OpenStudy (anonymous):

Solve. 4y + 16 > –8 A. y > –6 B. y > –2 C. y > 2 D. y > 6

OpenStudy (freckles):

Subtract 16 on both sides. Then last step is to divide both sides by 4. All I'm doing is undoing the operations near y to solve for y. 4y+16>-8 I see then +16 so I say okay I will -16 on both sides I see the 4* so I say okay I will div 4 on both sides.

OpenStudy (anonymous):

-8 - 16 = -24 right

OpenStudy (freckles):

so you have 4y>-8-16 that means 4y>-24

OpenStudy (anonymous):

y+6

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

divide both sides of 4y>-24 by 4

OpenStudy (anonymous):

so is it - 6

OpenStudy (freckles):

well you need the inequality but yeah you have y inequality sign goes here -6

OpenStudy (anonymous):

so im right?

OpenStudy (freckles):

well as I said you are missing your inequality symbol

OpenStudy (freckles):

if you are asking about -24 divided by 4 only then yeah that is -6

OpenStudy (freckles):

but without the y (inequality sign) part the solution is incomplete

OpenStudy (anonymous):

so its a?

OpenStudy (freckles):

4y>-24 since we divided by a positive the inequality direction still holds

OpenStudy (freckles):

so you have y>-6